OpenStudy (anonymous):

Find the radius of convergence for the series from n=1 to infinity of (n^n * x^n) / n!

OpenStudy (amistre64):

hows your limit coming along?

OpenStudy (anonymous):

I am using the ratio test. I know the taylor series of e^x is series of x^n / n!

OpenStudy (anonymous):

The limit I got (n+1)^(n+1) * x / (n+1)

OpenStudy (anonymous):

Isn't the numerator always greater so it would be divergent by the divergent test?

OpenStudy (amistre64):

\[|x|\lim~\frac{n!~(n+1)^{n+1}}{(n+1)!~n^n}\] \[|x|\lim~\frac{(n+1)^{n+1}}{(n+1)~n^n}\] \[|x|\lim~\frac{(n+1)^{n}}{~n^n}\]

OpenStudy (anonymous):

We want |n+1/n| to be < 1

OpenStudy (amistre64):

not that i see, we just want the limit now

OpenStudy (anonymous):

the limit is 1 > 0 which diverges.

OpenStudy (anonymous):

but we still have |x| < 1

OpenStudy (amistre64):

no, as long as the limit is some finite value, its fine ... then we can work on teh radius of convergence

OpenStudy (anonymous):

the radius of convergence is 1.

OpenStudy (amistre64):

is that what you are finding?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

well, im not finding that

OpenStudy (anonymous):

Oh... could you show me what you got?

OpenStudy (anonymous):

I thought the limit is x and we want |x|<1 by the ratio test.

OpenStudy (amistre64):

i showed you how i got to here \[|x|\lim~\frac{(n+1)^{n}}{~n^n}\] \[|x|\lim~\left(\frac{n+1}{n}\right)^n\] \[|x|\lim~\left(1 + \frac{1}{n}\right)^n\] now spose the limit of the n parts is some finite value \[\lim~\left(1 + \frac{1}{n}\right)^n=L\] then we work the setup: \[L|x|<1\] \[|x|<\frac1L\] and thats our radius

OpenStudy (amistre64):

do you know the value of L?

OpenStudy (anonymous):

I think it was 1.

OpenStudy (amistre64):

its not

OpenStudy (anonymous):

lim of (n+1)^n / n^n ? is similar to n^n/ n^n which is 1?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

\[\lim_{n \to \infty}\left(1+\frac1n\right)^n\implies e \]

OpenStudy (anonymous):

How did you get that?

OpenStudy (amistre64):

this is something that you prolly should have already been exposed to in a prior problem. its a bit lengthy to describe and im a bit out of practice. but if you know anything about a compounding rate of interest....

OpenStudy (anonymous):

Also, is there any way we could solve this problem by using this info: the taylor series of e^x is series of x^n / n!

OpenStudy (anonymous):

Thank you so much btw :)

OpenStudy (amistre64):

im not sure if we can or not, the radius of convergence is not dependant on a 'parent' function f(x) e^x does not necessarily have the same radius of convergence as e^x

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

i wish i could recall a good way to prove the limits e, but nothings coming to mind :) been too long

OpenStudy (anonymous):

Alrighty, thank you for your help anyway :) I really appreciate it

OpenStudy (anonymous):

L'Hopital's rule will do the trick

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

Provided you write the limit in indeterminate form, of course

OpenStudy (anonymous):

sounds good, thank you:)